# Cross-Modal Learning for Anomaly Detection in Fused Magnesium Smelting Process: Methodology and Benchmark

### v1

@@ Abstract
Fused Magnesium Furnace (FMF) is a crucial industrial equipment in the production of magnesia, and anomaly detection plays a pivotal role in ensuring its efficient, stable, and secure operation. Existing anomaly detection methods primarily focus on analyzing dominant anomalies using the process variables (such as arc current) or constructing neural networks based on abnormal visual features, while overlooking the intrinsic correlation of cross-modal information. This paper proposes a cross-modal Transformer (dubbed FmFormer), designed to facilitate anomaly detection in fused magnesium smelting processes by exploring the correlation between visual features (video) and process variables (current). Our approach introduces a novel tokenization paradigm to effectively bridge the substantial dimensionality gap between the 3D video modality and the 1D current modality in a multiscale manner, enabling a hierarchical reconstruction of pixel-level anomaly detection. Subsequently, the FmFormer leverages self-attention to learn internal features within each modality and bidirectional cross-attention to capture correlations across modalities. By decoding the bidirectional correlation features, we obtain the final detection result and even locate the specific anomaly region. To validate the effectiveness of the proposed method, we also present a pioneering cross-modal benchmark of the fused magnesium smelting process, featuring synchronously acquired video and current data for over 2.2 million samples. Leveraging cross-modal learning, the proposed FmFormer achieves state-of-the-art performance in detecting anomalies, particularly under extreme interferences such as current fluctuations and visual occlusion caused by heavy water mist. The presented methodology and benchmark may be applicable to other industrial applications with some amendments. The benchmark will be released at https://github.com/GaochangWu/FMF-Benchmark.
@@ I Introduction
Fused magnesia possesses numerous properties such as high temperature resistance, strong oxidation resistance and corrosion resistance. It serves as an indispensable resource in clinical surgery, aerospace, industry and other fields [1]. Fused Magnesium Furnace (FMF) is the main equipment for the production of fused magnesia, which melts powdered ore by generating a molten pool around 2850C through electric arc. Since the entire production is a continuous alternation of ore feeding and smelting, coupled with dynamic fluctuations in ore properties and non-optimal settings of smelting current [2], the temperature of molten pool can become unstable, leading to abnormal conditions in the FMF. Typical anomalies, especially the semi-molten condition [3], are detrimental to product quality and even threaten production security if not resolved in time. Therefore, timely and accurate anomaly detection is essential to the high quality, stable, and secure production of the fused magnesium smelting process.
@@
Figure 1: Cross-modal information is exploited to detect abnormal condition in fused magnesium smelting processes, as illustrated in (a). The picture at the bottom left shows an anomaly region on the furnace shell, whose visual feature is difficult to detect due to interference from heavy water mist. A novel FMF Transformer (FmFormer) is proposed using synchronous acquired video and current data, to explore the internal features of each modality by self-attention and the correlation feature across modalities by cross-attention, as shown in (b).
@@
With the success of deep learning in artificial intelligence [6, 7], recent researches are stepping towards deep learning-based anomaly detection using image input [8, 9] or video [10, 11] acquired from industrial cameras. Comparing with smelting current, abnormal conditions of FMF demonstrate prominent visual features. Taking the semi-molten condition as an example, in its initial phase, undesired variation in the molten pool temperature is accompanied by a red dot in the local area of the furnace shell [10]. Based on this visual feature, Wu et al. [10] separated the anomaly detection task into the spatial feature extraction from 3D video using a 2D Convolutional Neural Network (CNN), and the temporal feature extraction and prediction using a recurrent neural network, i.e., Long-Short Term Memory (LSTM) [12]. Recently, Liu et al. [11] further explored a 3D convolutional LSTM [13] to learn spatio-temporal video data simultaneously. Although the vision-based anomaly detection provides a scheme with higher accuracy and better prediction consistency than the current-based detection, there are still challenges in such complex industrial environment due to visual interference (e.g., water mist and light impact from furnace flame, as shown in the bottom left image of Fig. 1(a)), which diminish detection fidelity of vision-based approaches. Therefore, it is crucial to utilize both video and current information for a comprehensive anomaly detection in fused magnesium smelting processes.
@@
To enhance anomaly detection in fused magnesium smelting processes, we attempt, in this paper, feature exploration from both visual and current information through the cross-modal bridge. Inspired by the scaling success of Transformer models in natural language processing [14, 15] and computer vision [16, 17, 18, 19], we propose a novel cross-modal Transformer, dubbed FmFormer, to explore both spatial (visual) and temporal (visual and current) features, as illustrated in Fig. 1(b). First, we present a novel multiscale tokenization paradigm that effectively samples 3D patches with varying local receptive fields from visual modality (video) and 1D vectors from current modality, and embeds the samples into features (also known as tokens). Despite the considerably large dimensionality gap between 3D video and 1D current data, the tokenization paradigm seamlessly converts the multi-modal inputs into features with equivalent dimensions, enabling efficient exploration of intrinsically correlated features using succeeding attention mechanisms. More importantly, the proposed multiscale tokenization paradigm facilitates pixel-level anomaly detection, progressing from coarse to fine through hierarchical reassembly of the multiscale tokens. After tokenization, the proposed FmFormer processes the internal features of each modality separately through a self-attention. Subsequently, a bidirectional cross-attention is utilized to obtain correlation features in both current-to-visual direction and visual-to-current direction. By leveraging the cross-modal learning, high-fidelity anomaly detection can be achieved by decoding correlated tokens for class-level predictions or assembling current-to-visual tokens to locate anomaly regions (pixel-level predictions).
@@
  • A novel cross-modal Transformer dubbed FmFormer, for both class-level and pixel-level anomaly detection in fused magnesium smelting processes. The FmFormer progressively encodes the internal features of each modality and the correlation feature across the modalities through a cascading structure of alternatively stacked self-attention and cross-attention.
  • A novel multiscale tokenization paradigm for generating token sets with varying local receptive fields to facilitate the hierarchical reconstruction of pixel-level anomaly predictions with high localization accuracy.
@@
The rest of the paper is organized as follows. Section II reviews previous works related to the proposed method, including anomaly detection and cross-modal learning. Section III provides a detailed description of our proposed FmFormer. Section IV introduces the proposed cross-modal benchmark, the training recipe, and the implementation details for the FmFormer. Finally, we evaluate our method on the presented cross-modal benchmark and perform ablation studies of the proposed modules in Section V. Section VI provides conclusion and limitations of this work.
@@ II-A Learning-based Anomaly Detection
Anomaly detection is defined as discovering patterns in data that do not match desired behavior, which can be mainly categorized into learning-based, statistical-based, information theory-based and spectral theory-based [20]. In this paper, we only focus on learning-based anomaly detection approaches. For time-series inputs, Yin et al. [21] converted the one-dimensional data into two-dimensional data using a sliding window scheme and implemented anomaly detection by combining convolutional layers and LSTM cells into an autoencoder architecture. Barz et al. [22] introduced a maximizes divergent intervals framework for spatio-temporal anomaly detection in an unsupervised learning manner. The proposed framework detects suspicious blocks generated by a different mechanism by modeling the probability density of data sub-blocks. To achieve unsupervised learning from only normal data, Liu et al. [23] introduced generative adversarial networks that learn to reconstruct the time-series signals in a low-dimensional space. When the reconstruction error is large, the input signal is considered to be abnormal. Due to the effective reduction of annotation cost, this unsupervised learning style has also been extended to image and video anomaly detection. However, significant noise and occlusion interference in harsh industrial environments greatly degrade model performance. For video anomaly detection with supervised learning, Zaheer et al. [24] introduced a generative adversarial learning method using 3D convolutional backbones. But convolution operation has a limited receptive field, resulting in biased local information.
@@
To solve the local-bias problem mentioned above, the self-attention mechanism in Transformer model [14] is designed to compute the correlation of each element to all the other elements, resulting in a global perception. Based on this structure, Xu et al. [25] introduced an anomaly Transformer model that is able to uniformly model local and global information of time-series data, which is challenging for a convolutional architecture. Chen et al. [26] proposed a bidirectional spatio-temporal Transformer to predict urban traffic flow using graph-based traffic representation. Cross-attentions of past-to-present and future-to-present directions are designed to learn the temporal tendency from the traffic data. To solve high-dimensional vision tasks, Dosovitskiy et al. [16] extended Transformer by decomposing an image into a sequence of vectorized patches and converting them into token representations. This extended version is called ViT. For visual anomaly detection using 3D video input, Li et al. [27] applied a convolutional Transformer to encode 2D image slices into feature vectors, and then used another convolutional Transformer to decode these vectors into detection result.
@@
Recently, a growing number of studies have shown that pure Transformer structures are also capable of learning representative features from high-dimensional video without using convolution-based encoders. For instance, Arnab et al. [28] extended the 2D tokenization of ViT [16] into 3D space via 3D convolution, called tubelet embedding. 3D patches are extracted from the input video to construct vectorized tokens, which are further fed into a standard Transformer. Piergiovanni et al. [29] generalized this idea and proposed to utilize tubes of different shapes to sparsely sample the input video. Since 3D patches and 2D patches are jointly extracted in the tokenization, both 3D video datasets and large-scale of 2D image datasets can be seamlessly applied for network training. Different from the aforementioned tokenization paradigms that samples 3D patches from the input video, we highlight the idea of dilated convolution [30] and propose a novel dilated tokenization to construct tokens with a larger local receptive field. This tokenization paradigm naturally forms a multiscale mechanism for efficiently reconstructing pixel-level anomaly detection.
@@ II-B Cross-Modal Learning
Human-beings inherently have the ability to perceive and process cross-modal information such as language, sound, image, etc. With the rapid increase in the diversity of acquired information and the improvement of computing power, cross-modal learning is becoming an emerging direction in the field of artificial intelligence, e.g., hybrid imaging [31, 32, 33, 34], visual question answering [35, 36, 37], visual-text retrieval [38, 39, 40, 41], robot perception [42, 43] and chatbot (the ChatGPT [44, 15]).
@@
A straightforward multi-modal learning scheme is to process each modality input with a different branch of network and then merge them to generate the fused feature. For instance, Zhou et al. [8] introduced a multi-source information fusion method that applies a CNN-based image recognition branch and a current processing branch. In their study, the detection results from the multi-source are individually predicted by each branch and then fused together with a support vector machine. Bu et al. [45] further applied a CNN and a Multi-Layer Perceptron (MLP) to extract the image feature and current feature, respectively, and then used these features to predict anomalies. However, the above linear fusion scheme neglects the correlation between the modalities, resulting in the underutilization of multi-modal information. In comparison, the most significant difference between the proposed FmFormer and the aforementioned studies is that we explicitly model the information interaction between video and current by designing the cross-attention to explore the bidirectional correlation between these two modalities, i.e., cross-modal learning [46].
@@
Figure 2: An overview of the proposed FmFormer for anomaly detection in fused magnesium smelting processes. (a) The tokenization module converts 3D video and 1D current into two set of vectors, called tokens, embedded with positional information. We propose a novel multiscale video tokenization paradigm based on dilated sampling to obtain tokens with different spatial scales while keeping them with the same dimension. (b) The cross-modal Transformer encoder learns cross-modal representation from video tokens and current tokens by exploring internal features with self-attention and correlation features with bidirectional cross-attention. The bidirectional cross-attention integrates the features from the two modalities, generating current-to-visual tokens and visual-to-current tokens. (c) A multi-head decoder is developed to perform dense prediction and class prediction of abnormal conditions. The dense prediction head serves to spot anomaly regions by reassemble multi-scale video tokens (in light green) into image-like feature representations. The classification head fuses the class tokens from video and current inputs (in dark green and dark violet) into a classification vector.
@@
Explicit cross-modal learning involves modeling the correlation between two modalities in the embedding space. For example, Ben-younes et al. [35] model the visual question answering task as a bilinear interaction between visual and linguistic features, and introduced a Tucker decomposition of the correlation tensor to explicitly control the model complexity. To reduce the dimensionality gap between 3D video and 1D text, Lei et al. [47] employed a ResNet [48] to encode 3D video into a set of 2D features and squeezed the temporal dimension of visual features via an average-pooling. Video tokens and text tokens are then concatenated and fed into a Transformer encoder. Instead of compressing temporal features, Yang et al. [49] proposed to embed each video frame into a visual token via a well-pre-trained visual-language model [50]. Wu et al. [51, 52] revisited the video classification task and converted it to a cross-modal learning problem by constructing a correlation matrix for video and language embeddings. This video classification paradigm is able to leverage well-pre-trained language models to generate precise semantic information. Although the spatio-temporal correlation is explored in these studies, each sample (e.g., video frame) in the time dimension is treated equally. Wu et al. [37] further introduced a temporal concept spotting mechanism using video and language embeddings to produce a category dependent temporal saliency map, which helps to enhance video recognition. Following these researches, we explicitly learn the correlation between different modalities through cross-attention mechanism in the embedding space. The main difference is that we employ a cascade structure with alternatively stacked self-attention and cross-attention, where the former is used for internal feature exploration for each modality.

### v2

# Cross-Modal Learning for Anomaly Detection in Complex Industrial Process: Methodology and Benchmark

@@ Abstract
Anomaly detection in complex industrial processes plays a pivotal role in ensuring efficient, stable, and secure operation. Existing anomaly detection methods primarily focus on analyzing dominant anomalies using the process variables (such as arc current) or constructing neural networks based on abnormal visual features, while overlooking the intrinsic correlation of cross-modal information. This paper proposes a cross-modal Transformer (dubbed FmFormer), designed to facilitate anomaly detection by exploring the correlation between visual features (video) and process variables (current) in the context of the fused magnesium smelting process. Our approach introduces a novel tokenization paradigm to effectively bridge the substantial dimensionality gap between the 3D video modality and the 1D current modality in a multiscale manner, enabling a hierarchical reconstruction of pixel-level anomaly detection. Subsequently, the FmFormer leverages self-attention to learn internal features within each modality and bidirectional cross-attention to capture correlations across modalities. By decoding the bidirectional correlation features, we obtain the final detection result and even locate the specific anomaly region. To validate the effectiveness of the proposed method, we also present a pioneering cross-modal benchmark of the fused magnesium smelting process, featuring synchronously acquired video and current data for over 2.2 million samples. Leveraging cross-modal learning, the proposed FmFormer achieves state-of-the-art performance in detecting anomalies, particularly under extreme interferences such as current fluctuations and visual occlusion caused by heavy water mist. The presented methodology and benchmark may be applicable to other industrial applications with some amendments. The benchmark will be released at https://github.com/GaochangWu/FMF-Benchmark.
@@ I Introduction
Fused magnesium smelting is a typical complex industrial process that produces fused magnesia, which has numerous properties, including high temperature resistance, strong oxidation resistance, and corrosion resistance. These qualities make it an indispensable resource in fields such as clinical surgery, aerospace, and various industries [1]. Fused Magnesium Furnace (FMF) is the main equipment in the fused magnesium smelting process to melt powdered ore by generating a molten pool around 2850C through electric arc. Since the entire production is a continuous alternation of ore feeding and smelting, coupled with dynamic fluctuations in ore properties and non-optimal settings of smelting current [2], the temperature of molten pool can become unstable, leading to abnormal conditions in the FMF. Typical anomalies, especially the semi-molten condition [3], are detrimental to product quality and even threaten production security if not resolved in time. Therefore, timely and accurate anomaly detection is essential to the high quality, stable, and secure production of the fused magnesium smelting process.
@@
Figure 1: Cross-modal information is exploited to perform anomaly detection in the context of a typical industrial process, fused magnesium smelting, as illustrated in (a). The picture at the bottom left shows an anomaly region on the furnace shell, whose visual feature is difficult to detect due to interference from heavy water mist. A novel FMF Transformer (FmFormer) is proposed using synchronous acquired video and current data, to explore the internal features of each modality by self-attention and the correlation feature across modalities by cross-attention, as shown in (b).
@@
With the success of deep learning in artificial intelligence [6, 7], recent researches are stepping towards deep learning-based anomaly detection using image input [8, 9, 10] or video [11, 12] acquired from industrial cameras. Comparing with smelting current, abnormal conditions of FMF demonstrate prominent visual features. Taking the semi-molten condition as an example, in its initial phase, undesired variation in the molten pool temperature is accompanied by a red dot in the local area of the furnace shell [11]. Based on this visual feature, Wu et al. [11] separated the anomaly detection task into the spatial feature extraction from 3D video using a 2D Convolutional Neural Network (CNN), and the temporal feature extraction and prediction using a Recurrent Neural Network (RNN). Recently, Liu et al. [12] further explored a 3D convolutional Long-Short Term Memory (LSTM) [13] to learn spatio-temporal video data simultaneously. Although the vision-based anomaly detection provides a scheme with higher accuracy and better prediction consistency than the current-based detection, there are still challenges in such complex industrial environment due to visual interference. For instance, the detection accuracy can quickly degenerate when there is water mist or light impact from the furnace flame (as shown in the bottom left image of Fig. 1(a)), which occasionally occur in such complex industrial scenario. Given these limitations of single-modality information, it is crucial to utilize both video and current information for a comprehensive anomaly detection in fused magnesium smelting processes.
@@
To enhance anomaly detection in the context of the fused magnesium smelting process, we attempt, in this paper, feature exploration from both visual and current information through the cross-modal bridge. Inspired by the scaling success of Transformer models in natural language processing [14, 15] and computer vision [16, 17, 18, 19], we propose a novel cross-modal Transformer, dubbed FmFormer, to explore both spatial (visual) and temporal (visual and current) features, as illustrated in Fig. 1(b). First, we present a novel multiscale tokenization paradigm that effectively samples 3D patches with varying local receptive fields from visual modality (video) and 1D vectors from current modality, and embeds the samples into features (also known as tokens). Despite the considerably large dimensionality gap between 3D video and 1D current data, the tokenization paradigm seamlessly converts the multi-modal inputs into features with equivalent dimensions, enabling efficient exploration of intrinsically correlated features using succeeding attention mechanisms. More importantly, the proposed multiscale tokenization paradigm facilitates pixel-level anomaly detection, progressing from coarse to fine through hierarchical reassembly of the multiscale tokens. After tokenization, the proposed FmFormer processes the internal features of each modality separately through a self-attention. Subsequently, a bidirectional cross-attention is utilized to obtain correlation features in both current-to-visual direction and visual-to-current direction. By leveraging cross-modal learning, a multi-head decoder is tailored to process correlated tokens for class-level predictions and to assemble current-to-visual tokens for locating anomaly regions (pixel-level predictions), achieving accurate and robust anomaly detection.
@@
  • A novel cross-modal Transformer (FmFormer) is proposed with a cascading structure of alternately stacked self-attention and bidirectional cross-attention layers. It progressively encodes internal features of each modality and correlation features across modalities, fully leveraging the strengths of different modalities to achieve accurate and robust anomaly detection in the context of the fused magnesium smelting process.
  • A novel multiscale tokenization paradigm is tailored for generating token sets with varying local receptive fields. This paradigm facilitates the hierarchical reconstruction of pixel-level anomaly predictions with high localization accuracy.
  • A multi-head decoder is designed to translate correlated tokens into class-level and pixel-level predictions, enabling concurrent anomaly detection and anomaly localization.
@@ II-A Learning-based Anomaly Detection
Anomaly detection is defined as discovering patterns in data that do not match desired behavior, which can be mainly categorized into learning-based, statistical-based, information theory-based and spectral theory-based [20]. In this paper, we only focus on learning-based anomaly detection approaches. For time-series inputs, Yin et al. [21] converted the one-dimensional data into two-dimensional data using a sliding window scheme and implemented anomaly detection by combining convolutional layers and LSTM cells into an autoencoder architecture. Barz et al. [22] introduced a maximizes divergent intervals framework for spatio-temporal anomaly detection in an unsupervised learning manner. To achieve unsupervised learning from only normal data, Liu et al. [23] introduced generative adversarial networks that learn to reconstruct the time-series signals in a low-dimensional space. When the reconstruction error is large, the input signal is considered to be abnormal. Due to the effective reduction of annotation cost, this unsupervised learning style has also been extended to image and video anomaly detection. However, significant noise and occlusion interference in harsh industrial environments greatly degrade model performance. For video anomaly detection with supervised learning, Zaheer et al. [24] introduced a generative adversarial learning method using 3D convolutional backbones. But convolution operation has a limited receptive field, resulting in biased local information.
@@
To solve the local-bias problem mentioned above, the self-attention mechanism in Transformer [14] is designed to compute the correlation of each element to all the other elements, resulting in a global perception. Based on this structure, Xu et al. [25] introduced an anomaly Transformer that uniformly models both local and global information, which is challenging for convolution operations. Chen et al. [26] proposed a bidirectional spatio-temporal Transformer to predict urban traffic flow using graph-based traffic representation. Cross-attentions of past-to-present and future-to-present directions are designed to learn the temporal tendency from the traffic data. To solve high-dimensional vision tasks, Dosovitskiy et al. [16] extended Transformer by decomposing an image into a sequence of vectorized patches and converting them into token representations. This extended version is called ViT. For visual anomaly detection using 3D video input, Li et al. [27] applied a convolutional Transformer to encode 2D image slices into feature vectors, and then used another convolutional Transformer to decode these vectors into detection result.
@@
Recently, a growing number of studies have shown that pure Transformer structures are also capable of learning representative features from high-dimensional video without using convolution-based encoders. For instance, Arnab et al. [28] extended the 2D tokenization of ViT [16] into 3D space via 3D convolution, called tubelet embedding. 3D patches are extracted from the input video to construct vectorized tokens, which are further fed into a standard Transformer. Piergiovanni et al. [29] generalized this idea and proposed to utilize tubes of different shapes to sparsely sample the input video. Since 3D patches and 2D patches are jointly extracted in the tokenization, both 3D videos and 2D images can be seamlessly applied for network training. Different from the aforementioned tokenization paradigms that samples 3D patches from the input video, we highlight the idea of dilated convolution [30] and propose a novel dilated tokenization to construct tokens with a larger local receptive field. This tokenization paradigm naturally forms a multiscale mechanism for efficiently reconstructing pixel-level anomaly detection.
Despite the powerful modeling capabilities of Transformers, models using unimodal input [11, 9, 12, 10] in complex industrial environments remain limited. This limitation motivates us to combine both current and visual modalities for anomaly detection, leveraging their complementary strengths to enhance detection accuracy and robustness.
@@ II-B Cross-Modal Learning
Human-beings inherently have the ability to perceive and process cross-modal information such as language, sound, image, etc. With the rapid increase in the diversity of acquired information and the improvement of computing power, cross-modal learning is becoming an emerging direction in the field of artificial intelligence, e.g., hybrid imaging [31, 32, 33, 34, 35], visual question answering [36, 37, 38], visual-text retrieval [39, 40, 41], robot perception [42] and chatbot (the ChatGPT [43, 15]).
@@
A straightforward multi-modal learning scheme is to process each modality input with a different branch of network and then merge them to generate the fused feature. For example, Zhou et al.[8] introduced a multi-source information fusion method that employs a CNN-based image recognition branch and a current processing branch, with detection results individually predicted and then fused using a support vector machine. Similarly, Bu et al.[44] used a CNN and a Multi-Layer Perceptron (MLP) to extract image and current features, respectively, for anomaly prediction. However, the above linear fusion scheme neglects the correlation between the modalities, resulting in the underutilization of multi-modal information. The key difference between the proposed FmFormer and the aforementioned studies is that we explicitly model the information interaction between video and current through cross-attention, which explores the bidirectional correlation between these two modalities, i.e., cross-modal learning [45].
@@
Figure 2: An overview of the proposed FmFormer for anomaly detection in fused magnesium smelting processes: (a) The tokenization converts 3D video and 1D current into sets of vectors, or tokens, embedded with positional information. We introduce a novel multiscale video tokenization paradigm based on dilated sampling, which generates tokens of varying spatial scales while keeping consistent dimensions. (b) The cross-modal Transformer encoder learns representations from video and current tokens by exploring internal features with self-attention and correlation features through bidirectional cross-attention, producing current-to-visual and visual-to-current tokens. (c) A multi-head decoder is designed for dense and class prediction of abnormal conditions. The dense prediction head reassembles multi-scale video tokens (in light green) into image-like feature representations, while the classification head integrates class tokens from both modalities (in dark green and dark violet) into a classification vector.
@@
Explicit cross-modal learning involves modeling the correlation between two modalities in the embedding space. For example, Ben-younes et al. [36] model the visual question answering task as a bilinear interaction between visual and linguistic features, and introduced a Tucker decomposition of the correlation tensor to explicitly control the model complexity. Park et al. [35] developed spatial and channel Transformers to facilitate cross-modal learning and fusion of infrared and visible images with the same dimensions. To reduce the dimensionality gap between 3D video and 1D text, pre-trained CNN and RNN backbones are commonly used to align the dimensions across modalities, with modality interaction implemented either through cross-attention [39] or self-attention with concatenated features [46]. Instead of compressing visual features, Yang et al. [47] proposed to embed each video frame into a visual token via a well-pre-trained visual-language model [48]. Wu et al. [49, 50] revisited the video classification task and converted it to a cross-modal learning problem by constructing a correlation matrix for video and language embeddings. This video classification paradigm is able to leverage well-pre-trained language models to generate precise semantic information. Although the spatio-temporal correlation is explored in these studies, each sample (e.g., video frame) in the time dimension is treated equally. Wu et al. [38] further introduced a temporal concept spotting mechanism using video and language embeddings to produce a category dependent temporal saliency map, which helps to enhance video recognition.
Motivated by these studies, we employ bidirectional cross-attention to explicitly explore correlated features across modalities. The main differences between our FmFormer and existing cross-modal approaches are: i) We use a cascading structure to alternatively stack self-attention and cross-attention, enabling the network to progressively capture both internal features within each modality and correlations across modalities; ii) Our encoder processes visual tokens at multiple scales through a multiscale tokenization module, addressing the dimensionality gap between the modalities while enhancing robustness to visual interferences; iii) Our multi-head decoder reassembles multiscale visual tokens into image-like feature representations for hierarchical feature fusion, simultaneously tackling class-level and pixel-level prediction tasks for improved anomaly detection and localization accuracy.